### v1

@@ Abstract
This review, the third decadal evaluation of the nuclear physics of hydrogen-burning stars, is motivated by the great advances made in recent years by solar neutrino observatories, putting experimental knowledge of the proton-proton chain neutrino fluxes in the few-percent precision range. The basis of the review is a one-week community meeting held in July 2022 in Berkeley, California, and many subsequent digital meetings and exchanges.
Each of the relevant reactions of solar and quiescent stellar hydrogen burning is reviewed here, from both theoretical and experimental perspectives. Recommendations for the state of the art of the astrophysical S-factor and its uncertainty are formulated for each of them.
@@
I Introduction
The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place inside the Sun and other hydrogen-burning stars. It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
@@ I.1 Purpose
Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B]. This discrepancy became known as the Solar Neutrino Problem. An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B. Suggestions included plausible flaws in the SSM that might produce a cooler solar core, new particle physics such as neutrino oscillations, and changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies. By the mid-1990s, with new results from Kamioka II/III and the SAGE/GALLEX experiments indicating a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature, attention increasingly turned to neutrino oscillations and other potential particle physics solutions. Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
@@
At that time, during a meeting on the Solar Neutrino Problem hosted by the Institute for Nuclear Theory (INT), University of Washington, Seattle, a suggestion was made to convene the community working on solar nuclear reactions, in order to reach consensus on the best current values of cross sections and their uncertainties. The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics. The working groups that formed during this meeting worked over the next year to evaluate past work, determining the needed cross sections and their uncertainties. The results of this evaluation, which came to be known as Solar Fusion I or SF-I [1998RvMP...70.1265A], became the standard for use in solar modeling over the next decade.
@@
An update of SF I was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle. Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A]. Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle . An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [2006NuPhA.777....1A, 2009ARA&A..47..481A]. This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [2002RvMP...74.1073C]. The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”. While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H]. Reflecting the impact of new experimental work and improved theory, the SF II study of 2011RvMP...83..195A significantly revised the SF I cross section for the driving reaction of the CN (or CNO-I) cycle, and updated key pp-chain cross sections.
@@
The present SF-III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley. Sessions were held at the David Brower Center, Berkeley, and the Physics Department, UC Berkeley, with the assistance of N3AS staff. The workshop was co-hosted by N3AS, by Solar Fusion’s long-term sponsor, the INT, and by the European ChETEC-INFRA Starting Community for Nuclear Astrophysics222https://www.chetec-infra.eu/.
@@
The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field. Workshop participants were organized in nine working groups, who were charged with reviewing and evaluating work completed since SF II, to produce updated recommendations. The structure of this review largely follows that of the working groups. The timing of SF-III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied. SF-III will provide the best current nuclear astrophysics input for such modeling. This review summarizes that input, based on the recommendations of the nine SF-III working groups.
@@ I.2 Terminology used
where is the target nucleus, is the bombarding particle, and and are the reaction products, with distinguished as the heavier product. Frequently in nuclear astrophysics, the reaction cross section is re-expressed in terms of the factor , with related to as described below, and with the mass numbers of entrance channel nuclei and .
The relation between the nuclear cross section and astrophysical factor is
with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy. (In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with point Coulomb scattering. The quantity is known as the Sommerfeld parameter. Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
The rapid rise in the nuclear cross section with energy, reflecting the weakening of the Coulomb barrier due to the dependence of the Sommerfeld parameter, competes with a rapid decrease in the probability of finding interacting particles with the requisite center-of-mass energy , in the high-energy tail of the Maxwell-Boltzmann distribution for the solar plasma. Consequently, the thermonuclear reaction rate for a given stellar temperature takes the form
@@
For the solar reactions of interest, the energy of the Gamow peak is well below the height of the Coulomb barrier. The slowly varying can be expanded as a power series around , with the leading term being , and with corrections given by derivatives taken at ,
@@
For most of the nuclear reactions studied here, the notation is adopted, where and are the mass numbers of the projectile and target nuclei, respectively. A summary of SF-III reactions and recommended astrophysical factors are presented in Table 1.
@@
| – | – | LABEL:subsec:WG3:Total_$S$_factors | |||
@@ I.3 Scope and structure of the review
In the Sun, the primary mechanism for the conversion of four protons into 4He is the proton-proton chain of Fig. 1, during which two charge-changing weak interactions take place, each converting a proton into a neutron with the emission of a neutrino. The pp-chain consists of three main branches, ppI, ppII, and ppIII. These branches are distinguished by their dependence on the solar core temperature and by the neutrinos they produce, with the pp, 7Be, and 8B neutrinos serving as “tags” for the ppI, ppII, and ppIII chains, respectively. The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos will have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331]. By the mid-1990s solar neutrino experiments had established that the measured fluxes differed significantly from this expected pattern. This contributed to growing expectations that new neutrino physics might be the solution of the Solar Neutrino Problem. It also was an important motivation for SF I, as errors in the nuclear physics of the pp chains could distort the relationship between the various fluxes.
@@
The pp chains account for 99% of solar hydrogen burning and are the main focus of this review. The remaining 1% is generated through the CN cycle (or CNO-I cycle, see Fig. 2). Motivated in part by Borexino’s recent success in measuring the flux of solar CN neutrinos, this review also considers in more detail than past Solar Fusion reviews the reactions driving hydrogen burning in stars more massive than our Sun.
@@
The review is structured as follows. Section II begins with a summary of past solar neutrino experiments, highlighting some of the open questions that remain. It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that helped resolve the solar neutrino problem, as well as new detectors in various stages of development, including Hyper-Kamiokande, SNO+, JUNO, DUNE, and others.
@@
The cross section of the driving reaction of the pp chains, , is too small to be measured and thus must be taken from theory. This important reaction is reviewed in section III. All the other pp-chain reactions and decays have been studied in the laboratory, and used together with theory – which guides fits and extrapolations, and corrects for difference in the atomic environments of solar and terrestrial reactions – to predict solar fusion rates. The 2H(He, 3He(3He,2)4He, and 3HeB reactions are reviewed in Secs. IV, V, and VI, respectively. The minor hep pp-chain branch is reviewed in Section VII, and the electron-capture reactions on pp and 7Be are discussed in Section VIII. The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX
While a modest (%) contributor to solar energy production, the CN (or CNO-I) cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N. It is also the primary mode of hydrogen burning in stars more massive than the Sun, with central temperatures above 2K. This review treats the CNO cycles in somewhat more detail than in SF I and SF II. In SF II the cross section for the leading reaction, 14N()15O, was revised downward by a factor of two from the SF I recommended value. The SF-III value is close to that of SF II but with an updated uncertainty evaluation described in Section LABEL:sec:S114.
@@
A current focus of laboratory astrophysics is the higher-temperature pathways for hydrogen burning that operate in massive stars. Many of the contributing reactions have been re-measured. That progress is reviewed in Section LABEL:sec:OtherCNO. Reactions of the NeNa cycle are also treated in this section.
@@
Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by screening, the shielding of the nuclear charge by electrons. Because of their very different atomic environments, laboratory reactions and those occurring in the solar plasma are affected in distinct ways. Section LABEL:sec:Screening describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
@@
The energy transport inside the radiative zone is affected by the metals that were incorporated into the Sun when it first formed. While metals comprise less than 2% of the Sun by mass, they play an outsized role in determining the opacity. Very few of the needed radiative opacities have been measured in the laboratory, and the conditions under which measurements are made typically are not identical to those in the Sun. Section LABEL:sec:Opacities describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
@@
Much of the progress made in constraining the reactions of the pp-chain and CNO-I cycles has been enabled by new laboratories and the facilities they house. The underground LUNA laboratory is a prominent example [2018PrPNP..98...55B]. Section LABEL:sec:Facilities describes the capabilities of LUNA and other laboratories responsible for the measurements reviewed here.
II Solar neutrino observations
@@
Solar neutrinos offer a unique tool box for probing both the fundamental properties of these elusive particles, and their interactions with matter, as well as understanding their source: the fusion reactions that power our Sun. The original motivation for observations of solar neutrinos was precisely the hope to probe solar fusion. The first successful experiment beginning in 1967 at the Homestake mine [1998ApJ...496..505C] offered the surprising result of a neutrino flux suppressed to approximately one third of expectation. The intervening years have seen a tour de force of experimental efforts, leading to the resolution of the Solar Neutrino Problem and the confirmation of neutrino oscillations [1998PhRvL..81.1562F, 2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2003PhRvL..90b1802E, 2004PhRvL..92r1301A]. These high-precision flux and spectral measurements mapped out the details of solar fusion, probed the structure of the Sun, and demonstrated (together with atmospheric neutrino measurements) that neutrinos have masses and mix.

### v2

@@ Abstract
This review, the third decadal evaluation of the nuclear physics of hydrogen-burning stars, is motivated by the great advances made in recent years by solar neutrino observatories, putting experimental knowledge of the pp chain neutrino fluxes in the few-percent precision range. The basis of the review is a one-week community meeting held in July 2022 in Berkeley, California, and many subsequent digital meetings and exchanges.
The relevant reactions of solar and stellar hydrogen burning are reviewed here, from both theoretical and experimental perspectives. Recommendations for the state of the art of the astrophysical S-factor and its uncertainty are formulated for each of them.
@@
Introduction
The present review summarizes the state of our understanding, in the third decade of the 21st century, of the nuclear reactions and decays taking place in solar and stellar hydrogen burning. It also addresses related issues, including solar neutrino detection, radiative opacities, electron screening of nuclear reactions, and the status of the current facilities for measuring cross sections and opacities.
@@ I.1 Purpose
Ray Davis’s measurements of the flux of solar neutrinos [1968PhRvL..20.1205D, 2003RvMP...75..985D] showed a significant deficit with respect to the predictions of the standard solar model (SSM) [1966PhRvL..17..398B]. This discrepancy became known as the "solar neutrino problem." An intense debate ensued about possible explanations for this discrepancy, summarized in the entertaining review by 1988Natur.334..487B. Suggestions included plausible flaws in the SSM that might produce a cooler solar core (e.g. 1981A&A....96....1S; 1996PhRvL..77.4286C), new particle physics such as neutrino oscillations (1969PhLB...28..493G; 1978PhRvD..17.2369W), changes in the nuclear physics governing He synthesis, such as an unidentified low-energy resonance in the 3He+3He reaction affecting extrapolations of laboratory cross sections to solar energies [1972Natur.238...24F, 1972PhLB...40..602F], or the presence of dark matter particles modifying energy transport in the solar core [1985ApJ...294..663S]. By the mid-1990s, with new results from Kamioka II/III (1989PhRvL..63...16H; 1996PhRvL..77.1683F) and the SAGE/GALLEX (1991PhRvL..67.3332A; 1992PhLB..285..376A) experiments revealed a pattern of neutrino fluxes inconsistent with the expected scaling of those fluxes with the solar core temperature. Attention increasingly turned to neutrino oscillations and other potential particle physics solutions. Concerns about the nuclear physics of the SSM also evolved, focusing more on the uncertainties that might inhibit extraction of any such “new physics.”
@@
At that time, during a meeting on the Solar Neutrino Problem hosted by the Institute for Nuclear Theory (INT), University of Washington, Seattle, a suggestion was made to convene the community working on solar nuclear reactions, in order to reach consensus on the best current values of cross sections and their uncertainties. The INT hosted the proposed workshop in February, 1997, drawing representatives from almost every experimental group active in this area, as well as many of the theorists who were engaged in solar neutrino physics. The working groups that formed during this meeting worked over the following year to evaluate past work, determining the needed cross sections and their uncertainties. The results of this evaluation, which came to be known as Solar Fusion I or 1998RvMP...70.1265A [1998RvMP...70.1265A], became the standard for use in solar modeling over the following decade.
@@
An update of 1998RvMP...70.1265A was launched with the January, 2009, workshop “Solar Fusion Cross Sections II”, hosted again by the INT in Seattle. Initial results from SuperKamiokande, the Sudbury Neutrino Observatory, and Borexino were then in hand, and the conversion of approximately two thirds of solar electron neutrinos into other flavors had been firmly established [2002PhRvL..89a1301A]. Consequently, the motivation for the study was to ensure that solar model predictions would be based on the most current nuclear physics, so that meaningful uncertainties could be placed on neutrino parameters derived from solar neutrino measurements, such as the mixing angle . An additional source of uncertainty had arisen at the time, with the advent of three-dimensional (3D) radiative-hydrodynamic models of the Sun’s atmosphere for determining element abundances from photo-absorption lines [2006NuPhA.777....1A, 2009ARA&A..47..481A]. This improved analysis, though, lowered the inferred metallicities, leading to tension with the Sun’s interior sound speed profile determined from helioseismology [2002RvMP...74.1073C]. The discrepancy between higher metallicity models that accurately reproduce the Sun’s interior sound speed and lower metallicity models based on the most current treatment of the photosphere was termed the “solar composition problem”. While the neutrino fluxes were not known to the precision needed to distinguish between the competing models, it had been shown that the solar core metallicity could be extracted directly from future CN solar neutrino measurements, if the precision of associated nuclear cross sections were improved [2008ApJ...687..678H]. Reflecting the impact of new experimental work and improved theory, the 2011RvMP...83..195A study of 2011RvMP...83..195A significantly revised the 1998RvMP...70.1265A cross section for the driving reaction of the CNO-I cycle, and updated key pp-chain cross sections.
@@
The present SF III review began with a four-day workshop in July, 2022, “Solar Fusion Cross Sections III.” The meeting was hosted by Physics Frontier Center N3AS (Network for Neutrinos, Nuclear Astrophysics, and Symmetries)111https://n3as.berkeley.edu/, a consortium of institutions involved in multi-messenger astrophysics, whose central hub is the University of California, Berkeley. Sessions were held at the David Brower Center, Berkeley, and the Physics Department, UC Berkeley, with the assistance of N3AS staff. The workshop was organized by N3AS, by Solar Fusion’s long-term sponsor, the INT, and by the European ChETEC-INFRA Starting Community for Nuclear Astrophysics222https://www.chetec-infra.eu/.
@@
The 47 workshop participants represented most of the leading experimental and theoretical research groups that are active in the field. Workshop participants were organized into nine working groups, who were charged with reviewing and evaluating work completed since 2011RvMP...83..195A, to produce updated recommendations. The structure of this review largely follows that of the working groups. The timing of SF III was driven in part by planned large-scale asteroseismic surveys of hydrogen-burning stars such as ESA’s PLATO333https://platomission.com/ mission [2014ExA....38..249R], and the expectation that associated data analysis will require large libraries of stellar models in which masses, ages, metallicities, and other parameters are varied. SF III will provide the best current nuclear astrophysics input for such modeling. This review summarizes that input, based on the recommendations of the nine SF III working groups.
@@ I.2 Terminology used
where is the target nucleus, is the bombarding particle, and and are the reaction products, with distinguished as the heavier product. Frequently in nuclear astrophysics, the reaction cross section is re-expressed in terms of the S-factor , with S related to as described below, and with the mass numbers of entrance channel nuclei and .
The relation between the nuclear cross section and astrophysical S-factor is
with and the charge numbers of the target and projectile, the fine structure constant, the relative velocity of the interacting particles in units of , the reduced mass, and the center-of-mass energy. (In this review we use energy units of MeV unless otherwise stated.) This removes the sharp energy dependence from the cross section associated with s-wave Coulomb scattering off a point nucleus, thereby accounting for the leading penetration effects associated with the Coulomb barrier. The quantity is known as the Sommerfeld parameter. Consequently, unless there are resonances, varies gently, and at low energies can generally be expressed as a low-order polynomial in .
The S-factor in Eq. (2) is defined in terms of the nuclear (or “bare") cross section, while laboratory experiments are performed with targets consisting of neutral atoms. As electrons shield the nucleus, lowering the Coulomb barrier, their presence enhances the laboratory cross section. The S-factors derived from laboratory data must be corrected for this effect. As the energy is lowered, the screening correction becomes increasingly important. Thus particular care must be taken when extrapolating laboratory cross sections to lower energies, as Fig. 4 of 2011RvMP...83..195A illustrates. Unless otherwise noted, the S-factors discussed here are those for bare nuclei.
The rapid rise in the nuclear cross section with energy, reflecting the higher probability of penetrating the Coulomb barrier, competes with a rapid decrease in the probability of finding interacting particles with the requisite center-of-mass energy in the high-energy tail of the Maxwell-Boltzmann distribution. Consequently, the thermonuclear reaction rate for a given stellar temperature takes the form
@@
For the solar reactions of interest, the energy at the Gamow peak is well below the height of the Coulomb potential. The slowly varying can be expanded as a power series around , with the leading term being , and with corrections given by derivatives taken at ,
@@
For most of the nuclear reactions studied here, the notation is adopted, where and are the mass numbers of the projectile and target nuclei, respectively. A summary of SF III reactions and recommended astrophysical S-factors are presented in Table 1.
Finally, we provide a short list of abbreviations that are frequently used throughout several sections:
- ANC
asymptotic normalization coefficient.
- BBN
Big Bang nucleosynthesis.
- CME
common-mode error.
- EFT
effective field theory, and variants: chiral (EFT), pion-less (EFT), hybrid (EFT*), etc.
- LO, NLO, NNLO
leading order, next to (next to) leading order, etc.
- SSM
Standard Solar Model.
@@
| – | – | LABEL:subsec:WG3:Total_Sfactors | |||
@@ I.3 Scope and structure of the review
In the Sun, the primary mechanism for the conversion of four protons into 4He is the pp chain of Fig. 1, during which two charge-changing weak interactions take place, each converting a proton into a neutron with the emission of a neutrino. The pp chain consists of three main branches, ppI, ppII, and ppIII. These branches are distinguished by their dependence on the solar core temperature and by the neutrinos they produce, with the pp, 7Be, and 8B neutrinos serving as “tags” for the ppI, ppII, and ppIII chains, respectively. The SSM predicts that the fluxes of pp, 7Be, and 8B neutrinos have an approximate relative temperature scaling of 1:T:T, where Tc is the solar core temperature [annurev:/content/journals/10.1146/annurev.aa.33.090195.002331]. By the mid-1990s, solar neutrino experiments established that the measured fluxes differed significantly from this expected pattern. This contributed to growing expectations that new neutrino physics might be the solution of the solar neutrino problem. It also was an important motivation for 1998RvMP...70.1265A, as errors in the nuclear physics of the pp chain could distort the relationship between the various fluxes.
@@
The pp chain accounts for 99% of solar hydrogen burning and are the main focus of this review. The remaining 1% is generated through the CNO-I cycle, see Fig. 2). Motivated in part by Borexino’s recent success in measuring the flux of solar CN neutrinos, this review also considers in more detail than past Solar Fusion reviews the reactions driving hydrogen burning in stars more massive than our Sun.
@@
The review is structured as follows. Section II begins with a summary of past solar neutrino experiments, highlighting some of the open questions that remain. It discusses the second-generation experiments Super-Kamiokande, the Sudbury Neutrino Observatory, and Borexino that contributed in the resolution of the solar neutrino problem. It also describes Hyper-Kamiokande, SNO+, JUNO, and DUNE, new detectors in various stages of construction.
@@
The cross section of the driving reaction of the pp chain, , is too small to be measured and thus must be taken from theory. This important reaction is reviewed in section III. With the exception of the minor hep branch, all other reactions and decays have been studied in laboratories, with measurements then combined with theory to predict rates under solar conditions. The 2H(He, 3He(3He,2)4He, and 3HeB reactions are reviewed in Secs. IV, V, and VI, respectively. The hep reaction is reviewed in Section VII, and the electron-capture reactions on pp and 7Be are discussed in Section VIII. The 7Be(,B reaction responsible for the high energy neutrinos measured in Super-Kamiokande is treated in Section IX.
While a modest contributor to solar energy production, the CNO-I cycle is a potentially important probe of core metallicity as its hydrogen burning is catalyzed by the Sun’s primordial C and N (Section LABEL:sec:OtherCNO).
@@
A current focus of nuclear astrophysics laboratories is the higher-temperature pathways for hydrogen burning that operate in massive stars. Many of the contributing reactions have been re-measured. That progress is reviewed in Section LABEL:sec:OtherCNO.
@@
Charged-particle nuclear reactions occurring at energies below the Coulomb barrier are affected by electron screening, the shielding of the nuclear charge by electrons. Because of their very different atomic environments, laboratory reactions and those occurring in the solar plasma are affected in distinct ways. Section LABEL:sec:Screening describes the current status of efforts to account for these differences, when solar cross sections are extracted from laboratory measurements.
@@
The energy transport inside the radiative zone is affected by the metals444In this work, we follow the astrophysical terminology where ”metals” refer to all chemical elements except for the two lightest ones, hydrogen and helium. that were incorporated into the Sun when it first formed. While metals comprise less than 2% of the Sun by mass, they play an outsized role in determining the opacity. Very few of the needed radiative opacities have been measured, and the conditions under which measurements are made typically are not identical to those in the Sun. Section LABEL:sec:Opacities describes the current state of the art, and discusses connections between opacities and solar composition that have complicated efforts to resolve the solar metallicity problem.
@@
Much of the progress made in constraining the reactions of the pp chain and CNO-I cycles has been possible thanks to new laboratories and the facilities they host. The underground LUNA laboratory is a prominent example [2018PrPNP..98...55B]. Section LABEL:sec:Facilities describes the capabilities of selected laboratories involved in the measurements reviewed here.
Solar neutrino observations
@@
Solar neutrinos offer a unique tool box for probing both the fundamental properties of these elusive particles, and their interactions with matter, as well as understanding their source: the fusion reactions that power our Sun. The original motivation for observations of solar neutrinos was precisely the hope to probe solar fusion. The first successful experiment beginning in 1967 at the Homestake mine [1998ApJ...496..505C] offered the surprising result of a neutrino flux suppressed to approximately one third of expectation. The intervening years have seen a tour de force of experimental efforts, leading to the resolution of the Solar Neutrino Problem and the confirmation of neutrino oscillations [1998PhRvL..81.1562F, 2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2003PhRvL..90b1802E, 2004PhRvL..92r1301A]. These high-precision flux and spectral measurements mapped out the details of solar fusion, probed the structure of the Sun, and demonstrated (together with atmospheric neutrino measurements) that neutrinos have masses and mix. Table 2 lists the pp-chain and CNO cycle neutrino sources, their endpoint ( decay) or line (electron capture) energies , and predicted SSM fluxes.
| Source | Eν (MeV) | GS98 | AAGS21 | MB22p | units (cm-2s-1) |
| pp | 0.420 | 5.96 (0.6%) | 6.00 (0.6%) | 5.95 (0.6%) | 1010 |
| pep | 1.442 | 1.43 (1.1%) | 1.45 (1.1%) | 1.42 (1.3%) | 108 |
| hep | 18.77 | 7.95 (31%) | 8.16 (31%) | 7.93 (30%) | 103 |
| 7Be | 0.862(a) 0.384(b) | 4.85 (7.4%) | 4.52 (7.3%) | 4.88 (8.1%) | 109 |
| 8B | 15 | 5.03 (13%) | 4.31 (13%) | 5.07 (15%) | 106 |
| 13N | 1.198 | 2.80 (16%) | 2.22 (13%) | 3.10 (15%) | 108 |
| 15O | 1.732 | 2.07 (18%) | 1.58 (16%) | 2.30 (18%) | 108 |
| 17F | 1.738 | 5.35 (20%) | 3.40 (16%) | 4.70 (17%) | 106 |
Notes: (a) 90% and (b) 10% of the 7Be neutrino flux, respectively.